The salinity of seawater is mainly due to the presence of KCl. The normal freezing point of seawater is -2.2 oC.  Diluted seawater is used to culture fertilized eggs of a specific species of fish. The osmotic pressure of the fish eggs against pure water is 14 bars at 15 oC. Calculate the volume (in L) of seawater you need to make 2 liters of isotonic solution for the fish eggs (the culturing condition is 15 oC, 1 bar). 
Answer #1

Normal freezing paint of sea wader = -2.2°C The depressim in freezing point = Tº - Ty = 0-(-2.2) AT4 = 2.2°C AT = i ktm {ks f

answered by: ANURANJAN SARSAM